# EXIT: Context-Aware Extractive Compression for Enhancing Retrieval-Augmented Generation

### v2

@@ 5 Main Results
Table 1 summarizes our evaluation results across multiple datasets and compression strategies. With the 8B reader, EXIT demonstrates strong generalization: although trained solely on HQA, it effectively addresses both single-hop (NQ, TQA) and multi-hop (2WIKI) queries under out-of-domain conditions. Compared to all baseline methods, EXIT consistently improves EM scores—for instance, by 1.3 and 2.0 points on NQ and TQA, and by even larger margins of 2.5 and 8.1 points on HQA and 2WIKI, respectively. Notably, these accuracy gains come with an average latency of just 0.8s, substantially faster than abstractive compression approaches.
The benefits of EXIT become more pronounced at larger scales. Using the 70B reader, EXIT surpasses the accuracy of all competing methods, averaging a 3.7-point improvement in EM and a 3.3-point improvement in F1 over the uncompressed baseline. On HQA, it achieves a 3.3-point EM gain while maintaining an efficient 3.5s latency—faster than using uncompressed documents and still competitive with the previously fastest method, RECOMP-Extr, but with significantly higher accuracy. EXIT’s effectiveness and efficiency, especially with larger models, make it a practical solution for large-scale QA applications.
@@ 6 Analyses
We conduct a series of analyses examining EXIT’s robustness, classification performance, latency factors, and design choices under various configurations. Additional experimental results and analyses are provided in Appendix B.
@@ 6.1 Robustness Analysis
6.2 Classification Performance
To better understand the effectiveness of our context-aware relevance classifier, we report row-normalized confusion matrices for both in-domain (HQA) and out-of-domain (2WIKI) datasets, as shown in Figure 4. On HQA, the classifier displays a perfectly balanced ability to recognize both relevant (“Yes”) and irrelevant (“No”) sentences, achieving over 90% precision and recall in each category. While, on the 2WIKI dataset, the classifier exhibits a slight drop in recall for “Yes” sentences, it still performs strong classification ability with over 70% recall and 90% precision. These results confirm that the classifier performs robustly in its training domain and generalizes reasonably well to unseen queries, yet we leave narrowing this discrepancy as a valuable future research direction.
@@ 6.3 Understanding End-to-End Latency Factors
While previous work has primarily focused on minimizing token counts to reduce reading time, we emphasize the importance of considering end-to-end latency, including compression, for building an efficient RAG pipeline. We provide a breakdown of the total end-to-end latency into read time and compression time, along with an analysis of the average number of tokens in compressed documents, as shown in Figure 5. Although some methods achieve extreme token reduction—e.g., RECOMP-Abst (6.4%) and CompAct (10.2%)—their lengthy compression stages (1.46s and 7.99s, respectively) negate these gains, resulting in overall inference times that exceed the uncompressed baseline. By contrast, EXIT retains a moderate token ratio (28.3%) but completes compression rapidly (0.36s), bringing total latency to 0.79s, a 23.3% improvement over the 1.03s needed without compression. This analysis highlights the importance of balancing token reduction and compression inference latency to achieve actual efficiency in the RAG pipeline. Also, our proposed method, EXIT, aligns closely with these objectives, offering a practical approach to context compression for RAG.
@@ 6.4 Ablation Studies
Configuration EM $\uparrow$ F1 $\uparrow$ # token $\downarrow$ Ours (Pos + H-Neg + Neg) 31.6 42.6 195.1 Pos + H-Neg 30.0 41.3 286.8 Pos + Neg 29.8 40.9 404.6 w/o Adaptive Sentence Selection 29.4 40.7 91.0 w/o Passage as context 30.4 42.3 157.4
@@
Adaptive Sentence Selection. We compare our adaptive selection mechanism to fixed-length selection. Although fixed-length selection achieves the lowest token count (91.0), it reduces EM by 2.2 points and F1 by 1.9 points. This underscores the importance of adaptively selecting sentences based on the complexity of the query and the retrieved documents, rather than using a static cap.
Context-Aware Extraction. We assess the impact of incorporating full document context when evaluating each sentence’s relevance. Removing surrounding context saves 38 tokens but lowers EM by 1.2 points, indicating that broader contextual awareness is crucial for maintaining answer accuracy, even if it slightly increases token count.
These findings confirm that a balanced training data strategy enhances robustness, that adaptive sentence selection ensures efficiency, and that full document consideration preserves accuracy. The classification performances under each ablation setting are reported in Appendix B.5.
6.5 Impact of Compressor Model Size and Compression Strategy
@@
Model Size Considerations. Figure 6 presents an ablation study examining how different base models influence EM scores and total latency. Note that all models trained within EXIT achieve superior accuracy compared to uncompressed baselines and fast compression under 2 seconds. Specifically, Gemma-2B, our base classifier model, achieves a favorable balance between effectiveness and efficiency, delivering 31.6 EM points in 0.78s. When moving to the largest model, Gemma-7B, the highest accuracy (32.4 EM) is achieved, yet it also inflates latency to 1.50s, slightly exceeding the time of uncompressed documents. These results suggest that scaling up model parameters can improve performance but may also compromise latency benefits, emphasizing the flexibility of our proposed framework by selecting an appropriate model depending on the user requirement.
@@ 7 Conclusion
We present EXIT, an efficient context compression framework for RAG systems that leverages parallel processing and context-aware, adaptive sentence selection. Our experiments demonstrate that EXIT achieves superior performance across both single-hop and multi-hop QA tasks while maintaining practical inference speeds. Despite being trained only on HQA, EXIT shows a strong zero-shot generalization ability and proves effective across a wide range of open-source models of varying sizes. These results suggest that efficient parallel extraction with smaller models can outperform larger abstractive approaches, offering a practical solution for real-world RAG applications.
@@
Our current approach relies on explicit sentence-level annotations to train the classifier. While these annotations were obtained manually in our experiments, they could potentially be automated through alternative means, such as by GPT-4 supervision or signals derived from the reader itself. We have not yet explored these automated annotation strategies, but doing so remains a promising avenue for future work. Additionally, our study primarily focuses on a general-domain setting, leaving questions about the classifier’s performance in specialized domains unanswered. Investigating how well our approach generalizes to domain-specific or highly specialized corpora presents another valuable direction for future research. Lastly, we focus on a single-step RAG pipeline, where retrieval occurs only once, excluding more complex pipelines Shao et al. (2023); Trivedi et al. (2023); Khattab et al. (2022). However, our proposed framework, EXIT, is orthogonal to these approaches and can be seamlessly integrated by compressing the retrieved documents from each retrieval step.
@@
Appendix A More Implementation Details
@@
A.1 Training Configuration
We trained the compressor model using the following settings:
@@ A.1 Training Configuration
  • Quantization: 4-bit with float16
@@
A.2 Model Selection and Training Time
Model selection was based on validation loss. Training required approximately 90 hours on our cluster.
A.3 Data Processing
@@ A.5 Prompt Templates
Full prompt templates for compression and QA tasks are provided in Tables 4 and 5, respectively.
A.6 Reproducibility
We will release our codebase, including dataset preprocessing scripts, evaluation protocols, and model checkpoints, upon publication. All random seeds are set to 42 to facilitate reproducibility.
Appendix B Additional Experimental Results and Analyses
@@
B.1 Performance with Sparse Retrieval
@@
Compressor Type HQA 2WIKI EM $\uparrow$ F1 $\uparrow$ # Token $\downarrow$ EM $\uparrow$ F1 $\uparrow$ # Token $\downarrow$ Top-5 Documents Original Docs - 28.2 39.1 755.8 (100.0) 19.6 25.9 789.7 (100.0) RECOMP-Abst Abs. 27.8 39.2 63.0 (8.3) 25.0 30.6 55.7 (7.1) CompAct Abs. 30.6 41.2 76.9 (10.2) 19.6 29.1 71.0 (9.0) Refiner Abs. 30.8 42.3 84.0 (11.1) 19.6 27.8 62.9 (8.0) RECOMP-Extr Ext. 28.2 37.5 93.1 (12.3) 11.8 19.1 99.1 (12.5) LongLLMLingua Ext. 28.6 39.7 230.3 (30.5) 22.2 27.4 236.6 (30.0) EXIT (Ours) Ext. 33.4 44.5 177.8 (23.5) 24.4 29.1 138.2 (17.5) Top-20 Documents Original Docs - 31.2 42.5 3009.5 (100.0) 23.0 30.6 3132.5 (100.0) RECOMP-Abst Abs. 29.0 39.7 69.7 (2.3) 22.8 28.2 52.3 (1.7) CompAct Abs. 31.6 43.0 109.5 (3.6) 20.4 28.7 113.2 (3.6) Refiner Abs. 28.4 38.8 136.1 (4.5) 18.8 26.9 108.4 (3.5) RECOMP-Extr Ext. 28.4 37.0 93.5 (3.1) 11.2 19.3 96.7 (3.1) LongLLMLingua Ext. 31.0 40.7 558.0 (18.5) 23.6 28.3 593.7 (19.0) EXIT (Ours) Ext. 35.2 46.9 411.3 (13.7) 27.2 32.4 312.7 (10.0)
@@
B.2 Performance with Proprietary Model

### v3

@@ 5 Main Results
Table 1 summarizes our evaluation results across multiple datasets and compression methods. With the 8B reader, EXIT demonstrates strong generalization: although trained solely on HQA, it effectively addresses both single-hop (NQ, TQA) and multi-hop (2WIKI) queries under out-of-domain conditions. Compared to all baseline methods, EXIT consistently improves EM scores—for instance, by 1.3 and 2.0 points on NQ and TQA, and by even larger margins of 2.5 and 8.1 points on HQA and 2WIKI, respectively. Notably, these accuracy gains come with an average latency of just 0.8s, substantially faster than abstractive compression methods.
The benefits of EXIT become more pronounced at larger scales. Using the 70B reader, EXIT surpasses the accuracy of all competing methods, averaging a 3.7-point improvement in EM and a 3.3-point improvement in F1 over the uncompressed baseline. On HQA, it achieves a 3.3-point EM gain while maintaining an efficient 3.5s latency—faster than using uncompressed documents and still competitive with the previously fastest method, RECOMP-Extr, but with significantly higher accuracy. EXIT’s effectiveness and efficiency, especially with larger models, makes it a practical solution for large-scale QA applications.
@@ 6 Analyses
We conduct a series of analyses examining EXIT’s robustness, classification performance, latency factors, and design choices under various configurations. Additional experimental results and analyses are provided in Appendix C.
@@ 6.2 Understanding End-to-End Latency Factors
While previous work has primarily focused on minimizing token counts to reduce reading time, we argue that end-to-end latency—including both compression and generation time—offers a more practical measure of efficiency for RAG pipelines.
Figure 4 presents a breakdown of total latency into compression time and reading time, along with the average token count of the compressed context. Methods such as RECOMP-Abst (6.2% token ratio) and CompAct (10.6%) achieve strong compression but suffer from long compression times (1.46s and 7.99s, respectively), resulting in overall latency that exceeds the uncompressed baseline. By contrast, EXIT achieves a balanced 31.2% token ratio and fast compression (0.36s), reducing total latency to 0.79s—a 23.3% improvement over the uncompressed baseline (1.03s).
We acknowledge that EXIT may retain more content when the question is simple or the retrieved documents are of high quality. However, our goal is not to maximize compression alone, but to minimize total latency while preserving answer accuracy. Aggressive compression in such cases could introduce unnecessary overhead or risk omitting useful context. EXIT’s sentence-level classification balances this trade-off, allowing rapid, context-aware compression with consistent gains in both speed and accuracy, as shown in Table 1.
This analysis highlights that efficient compression is not solely about reducing tokens—it is about achieving the best trade-off between compression time and model input size. EXIT is designed with this principle in mind, making it a practical solution for real-world RAG applications.
6.3 Computational Complexity and Efficiency Trade-offs
Table 2: TFLOPs, latency, and relative speedup of compression methods. Latency is measured on a single A100-80GB GPU. Speedup is relative to the uncompressed baseline (Original Docs).
Method Comp. TFLOPs Read. TFLOPs Latency (s) Speedup
Original Docs 13.04 0.95 1.00×
RECOMP-Abst 0.54 1.60 2.31 0.41×
CompAct 18.35 2.05 12.86 0.07×
Refiner 15.03 2.51 6.78 0.14×
RECOMP-Extr 0.27 1.84 0.39 2.43×
LongLLMLingua 46.39 5.06 0.82 1.16×
EXIT (ours) 35.44 2.96 0.88 1.08×
To strengthen the justification of EXIT’s efficiency, we analyze not only latency but also computational complexity, measured in TFLOPs, using DeepSpeed’s FlopsProfiler Rasley et al. (2020). Table 2 reports the total FLOPs required for compression and reading (inference), as well as the resulting latency, all measured on a single A100-80GB GPU.
Although EXIT incurs higher total TFLOPs (38.40) than some competing methods, its design enables parallel sentence-level classification, which significantly reduces end-to-end latency. By contrast, abstractive methods like CompAct and Refiner involve autoregressive decoding, which must be executed sequentially and is subject to memory I/O bottlenecks, leading to much longer runtimes (up to 12.86 seconds).
This reveals a key trade-off: EXIT utilizes more GPU computation, but this computation is efficiently executed in parallel. Since many LLM inference pipelines are memory-bound rather than compute-bound, EXIT takes advantage of underutilized compute capacity to reduce latency. For example, compared to the uncompressed baseline (0.95s), EXIT achieves a lower latency of 0.88s, despite higher FLOPs.
These results demonstrate that EXIT provides practical efficiency by converting increased computation into latency reduction through parallel execution, offering a better balance between accuracy and response time than autoregressive alternatives.
@@ 6.4 Ablation Studies
Configuration EM $\uparrow$ F1 $\uparrow$ # token $\downarrow$ Ours (Pos + H-Neg + Neg) 31.6 42.6 195.1 Pos + H-Neg 30.0 41.3 286.8 Pos + Neg 29.8 40.9 404.6 w/o Adaptive Sentence Selection (2 sents) 29.4 40.7 91.0 w/o Adaptive Sentence Selection (4 sents) 30.2 41.4 166.5 w/o Passage as context 30.4 42.3 157.4
@@
Adaptive Sentence Selection. We compare our adaptive selection mechanism to a fixed-length strategy, which limits selection to four sentences. While the fixed-length method reduces token count to 166.5, it lowers EM by 1.4 and F1 by 1.2 points. This highlights the importance of adapting sentence selection based on query complexity and document relevance rather than imposing a static constraint.
Context-Aware Extraction. We evaluate the effect of considering full document context when determining sentence relevance. Removing surrounding context reduces token usage by 38 tokens but decreases EM by 1.2 points, highlighting the importance of contextual awareness in preserving answer accuracy. While context-aware extraction slightly increases token count, it ensures more precise and reliable responses.
These findings confirm that a balanced training data strategy enhances robustness, that adaptive sentence selection ensures efficiency, and that full document consideration preserves accuracy.
6.5 Compressor Model Size and Strategy Impact
@@
Model Size Considerations. Figure 5 presents an ablation study examining how different base models influence EM scores and total latency. Note that all models trained within EXIT achieve superior accuracy compared to uncompressed baselines and fast compression under 2 seconds. Specifically, Gemma-2B, our base classifier model, achieves a favorable balance between effectiveness and efficiency, delivering 31.6 EM points in 0.78s. When moving to the largest model, Gemma-7B, the highest accuracy (32.4 EM) is achieved, yet it also inflates latency to 1.50s, slightly exceeding the time of uncompressed documents. These results suggest that scaling up model parameters can improve performance but may also compromise latency benefits, emphasizing the flexibility of our proposed method by selecting an appropriate model depending on the user requirement.
@@
Table 4: Performance comparison across different retrieval and reranking methods on 2WIKI.
Method EM F1 #tok.
Passage-Level (Top-1)
bge-m3 14.6 22.7 156.5
bge-reranker-v2-gemma 15.0 22.9 156.2
RepLLaMA 16.0 24.1 157.2
UPR 9.2 16.8 155.3
Yes/No Ranking Prompt 13.4 21.8 157.5
Pairwise Ranking Prompt 13.2 21.5 156.2
Sentence-Level (Top-5)
bge-m3 16.6 24.2 216.9
bge-reranker-v2-gemma 18.4 26.2 203.2
RepLLaMA 14.8 23.2 239.0
UPR 7.4 14.6 148.9
Yes/No ranking prompt 14.6 21.8 135.6
Pairwise ranking prompt 11.6 19.4 209.9
Ours 24.8 30.1 150.2
6.6 Comparison with Reranking Methods
We compare EXIT with post-retrieval reranking methods in Table 4, including bge-m3 Chen et al. (2024a), bge-reranker-v2-gemma Li et al. (2023a), RepLLaMA Ma et al. (2024), UPR Sachan et al. (2022), Yes/No ranking prompting Liang et al. (2023), and Pairwise ranking prompting Qin et al. (2024). EXIT consistently achieves higher EM and F1 scores, demonstrating its ability to adaptively select query-relevant sentences while filtering out redundancy. By contrast, Top-$k$ re-ranking methods rely on static ranking, often retaining irrelevant content, increasing token count, and diminishing output quality. Moreover, they lack sentence-level context awareness, treating sentences independently without considering their broader document context, which can lead to incoherent or incomplete retrieval. These results underscore EXIT’s advantage in providing a more efficient and contextually relevant input for RAG than other post-retrieval methods.
6.7 Leveraging LLMs for Pseudo-Annotation
Figure 6: Confusion matrix of GPT-4o zero-shot classifier on HotpotQA. The model demonstrates high precision for identifying irrelevant sentences (“No”) but shows moderate recall for relevant ones (“Yes”).
To explore scalable alternatives to manual sentence-level annotations, we evaluated the use of a large language model (GPT-4o) as a zero-shot relevance classifier for generating pseudo-labels. Without any fine-tuning, GPT-4o achieved strong filtering precision, correctly identifying 97% of irrelevant sentences, though with lower recall (54%) for relevant ones (Figure 6). Despite this, EXIT trained with GPT-4o annotations still delivered competitive QA performance (Table 14), highlighting the feasibility of LLM-based pseudo-annotation as a scalable supervision strategy. These findings suggest a promising direction for improving generalization across domains by leveraging LLM-generated training data beyond curated datasets like HQA.
@@ 7 Conclusion
We present EXIT, an efficient context compression framework for RAG systems that leverages parallel processing and context-aware, adaptive sentence selection. Our experiments demonstrate that EXIT achieves superior performance across both single-hop and multi-hop QA tasks while maintaining practical inference speeds. Despite being trained only on HQA, EXIT shows a strong zero-shot generalization ability and proves effective across a wide range of open-source models of varying sizes. These results show that parallel extraction with smaller models outperforms larger abstractive methods, making it practical for real-world RAG.
@@
This work focuses on RAG pipelines, and its effectiveness in general long-context scenarios like LongBench Bai et al. (2024) remains future work. Our approach relies on explicit sentence-level annotations for training, which were manually obtained in our experiments. While these could be automated using GPT-4 supervision or reader-derived signals (as discussed in Section C.8), we have not yet explored fully automated annotation strategies—an important direction for future work.
EXIT performs sentence-level extraction, which may introduce limitations when sentences are overly long, noisy, or contain irrelevant details. While our context-aware classifier helps mitigate this, some ambiguity may persist, especially in cases involving complex or entity-heavy sentences. Additionally, although preserving the original sentence order empirically yields the best performance, we observed that removing intermediate sentences can occasionally lead to unnatural or incoherent reconstructions. Addressing this may require future improvements such as sentence decomposition, coherence-aware selection, or lightweight sentence refinement mechanisms.
We also focus on a single-step RAG setting, excluding iterative or recursive retrieval Shao et al. (2023); Trivedi et al. (2023); Khattab et al. (2022). However, EXIT is orthogonal to these approaches and can be integrated by compressing retrieved content at each step.
Finally, while EXIT avoids the latency bottlenecks of autoregressive generation, further efficiency gains could be achieved through architectural optimizations such as prefix caching or RadixAttention Zheng et al. (2024), which remain open areas for future exploration.
@@
Acknowledgments
This work was supported by the Institute for Information and communications Technology Promotion (IITP) grant funded by the Korea government (MSIT) (No. 2022-0-00010, Development of Korean sign language translation service technology for the deaf in medical environment).
@@ Acknowledgments
Appendix A More Related Work
A.1 Information Retrieval
Advancements in information retrieval (IR) have improved retrieval efficiency, granularity, and ranking quality. Encoder-based models, such as DPR Karpukhin et al. (2020) and Contriever Izacard et al. (2022), have been widely adopted for retrieval tasks. More recently, M3-Embedding Chen et al. (2024a) unifies dense, multi-vector, and sparse retrieval via self-knowledge distillation, enabling retrieval across different text granularities.
Decoder-only LLMs have emerged as strong rerankers, leveraging large-scale data for improved ranking at the cost of computational complexity. LLaRA Li et al. (2023a) and RePLlama Ma et al. (2024) employ dense retrieval and multi-stage ranking, while UPR Sachan et al. (2022) and ranking prompting methods Liang et al. (2023); Qin et al. (2024) refine retrieval through generation-based approaches. However, these methods focus on reordering passages rather than compressing retrieved content.
Research on retrieval granularity Seo et al. (2019); Lee et al. (2021); Jeong et al. (2023); Chen et al. (2024b); Hwang et al. (2024) explores balancing precision and contextual coherence. While fine-grained retrieval enhances specificity, excessive fragmentation can distort meaning Choi et al. (2021). Retrieval and ranking aim to maximize recall but do not address post-retrieval redundancy. EXIT operates beyond retrieval and ranking, focusing on adaptive post-retrieval context compression, which selectively filters content to optimize both efficiency and relevance in RAG.
A.2 Summarization
Summarization methods are broadly categorized as extractive and abstractive. Extractive summarization retains factual consistency by selecting key sentences but often suffers from redundancy and coherence issues Cheng and Lapata (2016); Narayan et al. (2018). Abstractive summarization improves fluency and readability but is computationally intensive and prone to hallucination See et al. (2017); Lewis et al. (2020a). Despite these challenges, its ability to generate concise outputs has driven interest in its application Song et al. (2024b); Lee et al. (2024); Song et al. (2024a).
Query-focused summarization refines document content based on relevance to a given query, removing unrelated information while preserving key details Zhang et al. (2024); Xu et al. (2023). However, traditional summarization methods process single documents and prioritize completeness and faithfulness Song et al. (2024a, b). By contrast, post-retrieval compression in RAG requires query-adaptive filtering across multiple retrieved documents. EXIT fills this gap by providing a dynamic and query-adaptive compression strategy, distinguishing between relevant, marginal, and irrelevant content while maintaining both efficiency over latency and answer accuracy.
Appendix B More Implementation Details
@@
B.1 Training Configuration
We employed LoRA Hu et al. (2022) to fine-tune the compressor model, enabling parameter-efficient training while preserving performance. The model was trained with the following hyperparameters:
@@ B.1 Training Configuration
  • Quantization: 4-bit with float16 precision
@@
B.2 Model Selection and Training Time
Model selection was guided by validation loss. Training was conducted on our cluster, requiring approximately 90 hours.
B.3 Data Processing
@@ B.5 Prompt Templates
Full prompt templates for compression and QA tasks are provided in Tables 6 and 7, respectively.
Appendix C Additional Experimental Results and Analyses
@@
C.1 Performance with Sparse Retrieval
@@
Compressor Type HQA 2WIKI EM $\uparrow$ F1 $\uparrow$ #token (%) $\downarrow$ EM $\uparrow$ F1 $\uparrow$ #token (%) $\downarrow$ Top-5 Documents Original Docs - 28.2 39.1 755.8 (100.0) 19.6 25.9 789.7 (100.0) RECOMP-Abst Abs. 27.8 39.2 63.0 (8.3) 25.0 30.6 55.7 (7.1) CompAct Abs. 30.6 41.2 76.9 (10.2) 19.6 29.1 71.0 (9.0) Refiner Abs. 30.8 42.3 84.0 (11.1) 19.6 27.8 62.9 (8.0) RECOMP-Extr Ext. 28.2 37.5 93.1 (12.3) 11.8 19.1 99.1 (12.5) LongLLMLingua Ext. 28.6 39.7 230.3 (30.5) 22.2 27.4 236.6 (30.0) EXIT (Ours) Ext. 33.4 44.5 177.8 (23.5) 24.4 29.1 138.2 (17.5) Top-20 Documents Original Docs - 31.2 42.5 3009.5 (100.0) 23.0 30.6 3132.5 (100.0) RECOMP-Abst Abs. 29.0 39.7 69.7 (2.3) 22.8 28.2 52.3 (1.7) CompAct Abs. 31.6 43.0 109.5 (3.6) 20.4 28.7 113.2 (3.6) Refiner Abs. 28.4 38.8 136.1 (4.5) 18.8 26.9 108.4 (3.5) RECOMP-Extr Ext. 28.4 37.0 93.5 (3.1) 11.2 19.3 96.7 (3.1) LongLLMLingua Ext. 31.0 40.7 558.0 (18.5) 23.6 28.3 593.7 (19.0) EXIT (Ours) Ext. 35.2 46.9 411.3 (13.7) 27.2 32.4 312.7 (10.0)
@@
C.2 Performance with Proprietary Model